# A new perspective on modulus stabilization

### v2

@@ IV.1 Flipping signs in the original GW setup
where $\gamma=\sqrt{4-\frac{m^{2}}{k^{2}}}$. Thus, to get a real scalar field (avoiding oscillatory solutions), we need to have $m^{2}\leq 4k^{2}$. This is well in agreement with [22] which says that in AdS space, the mass squared parameter $m^{2}$ for the GW scalar $\Phi$ can be negative without giving rise to instabilities as long as the condition $m^{2}/k^{2}+4>0$ is satisfied.
The intermediate equations (as in Sec. II(C)) get reproduced in the same manner except the $\nu$ getting replaced by $\gamma$ in all the expressions.
@@
$\begin{split}V_{\Phi}(r_{c})=-k\epsilon v_{h}^{2}+4ke^{-4kr_{c}\pi}(v_{v}-v_{h}e^{\epsilon kr_{c}\pi})^{2}\left(1-\frac{\epsilon}{4}\right)\\ +k\epsilon v_{h}e^{-(4-\epsilon)kr_{c}\pi}(2v_{v}-v_{h}e^{\epsilon kr_{c}\pi})\end{split}$ (28)
@@
Thus, we see that if $v_{v}$ and $v_{h}$ retain their previously assigned values (i.e. keeping the brane potentials fixed) which made the original GW scenario stable, we no longer have a modulus stabilization. The reason for this is that we had $v_{v}<v_{h}$ in the original setup which renders $kr_{c}$ to be negative (unphysical) from Eq.(29). What it essentially means is that we get a monotonic radion potential in the positive $r_{c}$ sector. Hence, we see that given fixed brane potentials, flipping sign of either term in the GW action results in modulus destabilization. An obvious question is - what if both the terms were made to flip signs? That would result in a violation of the weak energy condition (WEC) for the energy-momentum (EM) tensor of the scalar field (the original EM tensor for the scalar field was assumed to satisfy the WEC - so the new EM tensor which is just the negative of the previous no longer satisfies WEC [^2] WEC states that $T_{\mu\nu}u^{\mu}u^{\nu}\geq 0$ where $T_{\mu\nu}$ is the energy-momentum tensor and u is any timelike vector.). It is worth mentioning now that for each of the cases that we discuss here, when we discuss the inverted potential, we would be working with the same brane potentials and boundary conditions. This is essential to ensure a consistent comparison. In other words, we could have chosen the brane potentials such that we get a minimum in the inverted case but under the same conditions, we would not get an extremum in the original GW case.
@@
IV.2 Bulk cubic interaction
@@
Note that this solution is ill-defined at $kr_{c}\phi=-1/\xi$ which is an artifact of the analytic solution failing to be a good approximation. By appropriately choosing $\xi$, we can ensure $-1/\xi>>k\pi r_{c}$ and hence trust our solution. The radion potential coming from the GW part of the action is then given by
[…]
By plugging the solution for $\Phi$ (Eq.32) into the GW action and integrating over the extra dimension, we get the GW contribution to radion potential. Including the gravitational (GR) contribution and taking upto the leading order in $v$, the radion potential is given by [5]
$V(\varphi)=k^{4}\left(\frac{\varphi}{A}\right)^{4}\left[\tau+\frac{1}{1-\xi\log(\varphi/A)}\left(2\alpha v+\frac{\alpha^{2}\xi}{8}\right)\right]\;$ (34)
The potential is minimized at $\varphi\right>=f$ as
$\begin{split}\tau+\frac{w}{1-\xi\log(f/A)}=0\;\implies\left[\frac{f}{A}\right]^{\xi}\equiv e^{-\xi kr_{c}\pi}\\ =e^{1+w/\tau}+\mathcal{O}(\xi)\;\end{split}$ (35)
where $w=2\alpha v+\frac{\alpha^{2}\xi}{8}$. [^3] As $\tau=\Big{(}T_{vis}-\frac{\Lambda_{b}}{k}\Big{)}/k^{4}$ is present in the GR contribution to radion potential, we may fine-tune the visible brane tension accordingly to set the minimum of the net potential (GR+GW) at zero. In terms of the stabilized modulus, this translates to
$k\pi r_{c}\approx-\frac{\tau+w}{\xi\tau}\approx-\frac{1}{\xi}\;$ (36)
Thus, we see that the stabilized modulus depends directly on $\xi$ (thus $\eta$) and hence its signature. For a positive $\eta$, we get $r_{c}$ to be negative - hence, unphysical and no modulus stabilization (assuming $v$ to be positive). Thus, $\eta$ is chosen to be negative to ensure consistency, yielding a stabilized modulus. To obtain a large warp factor $k\pi r_{c}$ must be larger than one. This condition is satisfied if $v$ is of order $10^{-2}$ which validates the last approximation in Eq.(36).
We notice from Eq.(34) that if $\tau$ was neglected, the radion potential would be of the form $\frac{(2\alpha v+\frac{\alpha^{2}\xi}{8})}{1+\xi k\pi r_{c}}k^{4}e^{-4k\pi r_{c}}$. This, as we can see, goes to zero (its minimum) only when $r_{c}\to\infty$, unlike the previous case where we had a finite $r_{c}$. This would imply no modulus stabilization. This is especially interesting as it tells us something poignant about the underlying physics - the GW contribution alone cannot stabilize the modulus in the case of a bulk cubic interaction in the absence of gravitational contribution (as was in the original GW prescription). It is indeed the gravitational part of the radion potential which comes to the rescue. As a dominating bulk cubic interaction would imply an unbounded bulk potential for the GW scalar, this again provides evidence for a one-one correspondence between the GW scalar instability and modulus destabilization.
IV.3 Bulk quartic interaction
@@
where $\eta=\frac{kv^{2}\zeta}{12}$. The ill-definition of the solution as the denominator vanishes can be tackled in the same fashion as before. Now, plugging this in the GW action and integrating over the extra dimension gives us the radion potential:
[…]
@@
$\begin{split}\implies e^{-4kR}[\frac{k^{5}\eta^{2}v^{2}}{4}(1+\eta kR)^{-3}+\eta k^{5}v^{2}(1+\eta kR)^{-2}\\ -8\alpha k^{5}v(1+\eta kR)^{-1/2}-\alpha\eta vk^{5}(1+\eta kR)^{-3/2}+2\alpha^{2}k^{5}]=0\end{split}$ (42)
@@
$\begin{split}k\pi r_{c}[-\frac{3}{4}\eta^{3}v^{2}-2\eta^{2}v^{2}+4\alpha v\eta+\frac{3}{2}\alpha\eta^{2}v]=\alpha\eta v+8\alpha v\\ -\frac{\eta^{2}v^{2}}{4}-\eta v^{2}-2\alpha^{2}\end{split}$ (43)
@@
Thus, we arrive at an expression for the stabilized modulus for a bulk quartic scalar potential. It is indeed easy to see from Eq.(44) that if we were to work with an inverted quartic well ($\zeta\rightarrow-\zeta$), $r_{c}$ acquires a negative value which is unphysical, taking v to be positive. Hence, just like the standard GW case, we do not get modulus stabilization for the inverted case.
@@
IV.4 Bulk double-well potential
@@
We use Eq.(23) to find the OR solution for $\Phi$ which takes the form
@@
Figure 3: Effective radion potential for parameter values $b=0.2,v=0.01,k=10,\alpha=10^{-5}$ with $\lambda=0.001$ (left) and $\lambda=-0.001$ (right)
@@
What we get is that the effective potential has a global minimum for a double well bulk potential i.e. the modulus gets stabilized to a non-trivial non-zero value, which for our benchmark choice of parameters in Fig. 3 turns out to be 0.31. As a result, $k\pi r_{c}\approx 31$ which very well solves the hierarchy problem. However, for the inverted double well, the modulus gets destabilized (no minimum). Thus, we get another example which hints at a one-one correspondence between both the stabilities.
IV.5 Bazeia-Furtado-Gomes type bulk potential
Bazeia, Furtado and Gomes (BFG); in [23], introduced a set of bulk scalar potentials which are generated from a class of superpotentials of the form
@@ IV.5 Bazeia-Furtado-Gomes type bulk potential
$V_{p}(\Phi)=\frac{1}{8}\left(\frac{dW_{p}}{d\Phi}\right)^{2}-\frac{1}{3}W^{2}_{p}$ (52)
@@
The speciality of these potentials is that the coupled gravity-scalar system gives rise to thick brane solutions. What we do here is take such forms of bulk potentials and plug them in the RS-I background with proper brane potentials - hence solving for the scalar profile and backreacted metric using the superpotential approach of Ref. [4] (see Appendix A). The next step is to calculate the effective radion potential and check the behaviour for the entire class of superpotentials. A final question is if we can establish a relation between the value of the stabilized modulus ($r_{c}$) [if it gets stabilized] and the parameter p, which is addressed in Appendix B.
Following the steps as in Appendix A, we obtain the following solutions for the scalar profile and the warp factor [23]:
@@
$\begin{split}\noindent A_{p}(y)\!\!=\!\!-\frac{1}{3}\frac{p}{2p+1}\tanh^{2p}\left(\frac{y}{p}\right)-\frac{2}{3}\left(\frac{p^{2}}{2p-1}-\frac{p^{2}}{2p+1}\right)\\ \biggl{\{}\ln[\cosh\left(\frac{y}{p}\right)\biggr{]}-\sum_{n=1}^{p-1}\frac{1}{2n}\tanh^{2n}\left(\frac{y}{p}\right)\biggr{\}}\end{split}$ (54)
One should notice that, unlike all the cases till now, this is an exact solution involving the backreaction. The additional inputs that we need to give are the brane potentials which are precisely determined by Eqs. (48) and (49). Hence, putting in the scalar field solution and modified warp factor into the combined gravity-scalar action, integrating over the extra dimension and adding the contribution due to brane potentials; we get the effective radion potential  [^4] The Ricci scalar in the BFG convention is given by $R=8A"+20A^{\prime 2}$. This is incorporated in the numerics.. In this scenario as well, the calculations are quite cumbersome to do analytically, especially when our sole purpose is to determine the presence of modulus stabilization. So, we resort to a numerical approach.
@@
As the scalar field solution is a tan hyperbolic raised to an odd power, it has values between plus and minus one. This restricts the domain of the bulk potentials and hence, the bulk potentials are bounded from below by their values at ±$\pm 1$. Given the nature of our argument for all the cases till now, we would have naturally expected modulus stabilization in these scenarios which is indeed what we get.
IV.6 Mishra-Randall ansatz
Recently, Mishra and Randall [18] explored the effect of including bulk interaction terms for the GW scalar on RS cosmology. In particular, they were interested in the form of the beta function ($\beta=1/T$), where T is the
Figure 6: GW bulk potentials [($\epsilon_{2}<0,\epsilon_{3}<0$);($\epsilon_{2}<0,\epsilon_{3}>0$); ($\epsilon_{3}<0,\epsilon_{2}>0$); ($\epsilon_{2}>0,\epsilon_{3}>0$)]: from left to right
Figure 7: Radion potentials [($\epsilon_{2}<0,\epsilon_{3}<0$);($\epsilon_{2}<0,\epsilon_{3}>0$); ($\epsilon_{3}<0,\epsilon_{2}>0$); ($\epsilon_{2}>0,\epsilon_{3}>0$)]: from left to right
@@ IV.6 Mishra-Randall ansatz
$V_{hid}(\Phi)=\beta_{hid}(\Phi-v_{hid})^{2}\>,\;\>\beta_{hid}\to\infty\>,\\ \>V_{vis}(\Phi)=2\alpha_{vis}\Phi$ (56)
@@
with $0\leq r\leq r_{vis}$. Plugging everything into the GW action and integrating over the fifth dimension gives the following form for the radion potential [18]
$\begin{split}V(\varphi)=24M_{5}^{3}\kappa^{4}\,\varphi^{4}(1+\frac{a_{2}}{24M_{5}^{3}\kappa^{4}}\frac{\lambda\varphi^{\epsilon_{2}}}{1-\lambda\varphi^{\epsilon_{2}}}\\ -\frac{a_{3}}{24M_{5}^{3}\kappa^{4}}\,\log(1-\lambda\varphi^{\epsilon_{2}}))\end{split}$ (58)
@@
$\begin{split}\qquad\varphi=e^{-r_{vis}},\>\lambda=\frac{v_{\text{hid}}\epsilon_{32}}{1+v_{\text{hid}}\epsilon_{32}},\>\epsilon_{32}=\frac{\epsilon_{3}}{\epsilon_{2}},\\ \>a_{2}=-\frac{1}{32}\epsilon_{2}\alpha_{vis}^{2}-\frac{\epsilon_{2}}{\epsilon_{3}}\alpha_{vis}+2\alpha_{vis},\\ \>a_{3}=\frac{1}{2}\frac{\epsilon_{2}}{\epsilon_{3}}\alpha_{vis}\end{split}$ (59)
[…]
We get modulus stabilization for the $\epsilon_{3}=0,\epsilon_{2}<0$ case as is shown in Ref.[18]. However, the interesting case is $\epsilon_{2}=0$ i.e. only cubic contribution. We see from Eq.(59) that in this limit, $a_{2}=2\alpha_{vis},a_{3}=0,\epsilon_{32}\to\infty$ and hence, $\lambda\to 1$. Plugging these in Eq.(58), we see that the third term becomes zero while the second term is non-zero but blows up. The most important thing to notice here is that as all the $\phi$ contributions within the brackets were raised to the power $\epsilon_{2}$, we are left with no $\phi$ dependance for the quartic coefficient now - it is a constant. Thus, the radion VEV is either zero or driven to infinity depending on the sign of this constant implying no modulus stabilization irrespective of the sign of $\epsilon_{3}$. This is certainly what we expected as this was our conclusion when we independently examined the cubic bulk potential GW scenario without the gravitational contributions. Thus, it serves as a good check that our results match under proper limits.
V Addressing the correspondence
Here, we discuss two classes of the GW bulk potential (the potential satisfying certain properties) that clearly indicate the one-one correspondence between GW scalar stability and successful modulus stabilization. It is worthwhile to note that the below proofs include some approximations and hence, we would be critical and say that the correspondence is not exact - rather the possibilty of it being valid is incredibly high.
V.1 Global minimum for $V_{b}(\Phi)$

### v4

@@ 4.1 No bulk interactions
where $\gamma=\sqrt{4-\frac{m^{2}}{k^{2}}}$. Thus, to get a real scalar field (avoiding oscillatory solutions), we need to have $m^{2}\leq 4k^{2}$. This is well in agreement with the BF bound 1982 which says that in AdS space, the mass squared parameter $m^{2}$ for the GW scalar $\Phi$ can be negative without giving rise to instabilities as long as the condition $m^{2}/k^{2}+4>0$ is satisfied.
The intermediate equations (as in Ref.GW1) get reproduced in the same manner except the $\nu$ getting replaced by $\gamma$ in all the expressions.
@@
$\begin{split}V_{GW}(r_{c})=-k\epsilon v_{h}^{2}+4ke^{-4kr_{c}\pi}(v_{v}-v_{h}e^{\epsilon kr_{c}\pi})^{2}\left(1-\frac{\epsilon}{4}\right)+k\epsilon v_{h}e^{-(4-\epsilon)kr_{c}\pi}(2v_{v}-v_{h}e^{\epsilon kr_{c}\pi})\end{split}$ (24)
@@
Thus, we see that if $v_{v}$ and $v_{h}$ retain their previously assigned values (i.e. keeping the brane potentials fixed) which made the original GW scenario stable, we no longer have a modulus stabilization. The reason for this is that we had $v_{v}<v_{h}$ in the original setup which renders $kr_{c}$ to be negative (unphysical) from Eq.(25). What it essentially means is that we get a monotonic radion potential in the positive $r_{c}$ sector. Hence, we see that given fixed brane potentials, flipping sign of either term in the GW action results in modulus destabilization. An obvious question is - what if both the terms were made to flip signs? That would result in a violation of the weak energy condition (WEC) for the energy-momentum (EM) tensor of the scalar field (the original EM tensor for the scalar field was assumed to satisfy the WEC - so the new EM tensor which is just the negative of the previous no longer satisfies WEC [^2] WEC states that $T_{\mu\nu}u^{\mu}u^{\nu}\geq 0$ where $T_{\mu\nu}$ is the energy-momentum tensor and u is any timelike vector.). It is worth mentioning now that for each of the cases that we discuss here, when we discuss the inverted potential, we would be working with the same brane potentials and boundary conditions. This is essential to ensure a consistent comparison. In other words, we could have chosen the brane potentials such that we get a minimum in the inverted case but under the same conditions, we would not get an extremum in the original GW case. The equivalent analysis for CMS boundary conditions can be found in Sec.5.
@@
4.2 Bulk cubic interaction
@@
Note that this solution is ill-defined at $kr_{c}\phi=-1/\xi$ which is an artifact of the analytic solution failing to be a good approximation. By appropriately choosing $\xi$, we can ensure $1/\xi>>k\pi r_{c}$ and hence trust our solution. The radion potential coming from the GW part of the action is then given by
$\begin{split}V_{GW}(r_{c})=\int_{0}^{\pi}d\phi\>\frac{1}{r_{c}}e^{-4kr_{c}\phi}\Big(\partial_{\phi}\Phi\;\partial_{\phi}\Phi+\frac{r_{c}^{2}\,\eta}{3}\Phi^{3}\Big)+e^{-4kr_{c}\pi}2\alpha k^{5/2}\Phi(\pi)\;\end{split}$ (29)
By plugging the solution for $\Phi$ (Eq.28) into the GW action and integrating over the extra dimension, we get the GW contribution to radion potential.  [^3] Note $\varphi=Ae^{-k\pi r_{c}}$ is the canonical radion field. Chacko1
$V_{GW}(\varphi)=k^{4}\left(\frac{\varphi}{A}\right)^{4}\left[-\frac{\alpha^{2}}{4}+\frac{1}{1-\xi\log(\varphi/A)}\left(2\alpha v+\frac{\alpha^{2}\xi}{8}\right)\right]\;$ (30)
The potential is minimized at $\left<\varphi\right>=f$ as
$\begin{split}-\frac{\alpha^{2}}{4}+\frac{w}{1-\xi\log(f/A)}=0\;\implies\left[\frac{f}{A}\right]^{\xi}\equiv e^{-\xi kr_{c}\pi}\\ =e^{1-4w/\alpha^{2}}+\mathcal{O}(\xi)\;\end{split}$ (31)
where $w=2\alpha v+\frac{\alpha^{2}\xi}{8}$. [^4] The linear b.c. parameter $\alpha$ may be tuned accordingly to set the minimum of the GW radion potential at zero. In terms of the stabilized modulus, this translates to
$k\pi r_{c}\approx-\frac{1}{\xi}+\frac{4w}{\alpha^{2}\xi}\approx-\frac{1}{\xi}\;$ (32)
To obtain a large warp factor $k\pi r_{c}$ must be larger than one. This condition is satisfied if $v$ is of order $10^{-2}$ which validates the last approximation in Eq.(32). Thus, we see that the stabilized modulus depends directly on $\xi$ (thus $\eta$) and hence its signature. For a positive $\eta$, we get $r_{c}$ to be negative - hence, unphysical and no modulus stabilization (assuming $v$ to be positive).
4.3 Bulk quartic interaction
@@
where $\eta=\frac{kv^{2}\zeta}{12}$. The ill-definition of the solution as the denominator vanishes can be tackled in the same fashion as before. Now, plugging this in the GW action and dimensional reduction gives us the radion potential:
$\begin{split}V_{GW}(r_{c})=\int_{0}^{\pi}d\phi\>\frac{1}{r_{c}}e^{-4kr_{c}\phi}\Big(\partial_{\phi}\Phi\;\partial_{\phi}\Phi+\frac{r_{c}^{2}\,\zeta}{12}\Phi^{4}\Big)+e^{-4kr_{c}\pi}2\alpha k^{5/2}\Phi(\pi)\;\end{split}$ (35)
@@
$\begin{split}\implies e^{-4kR}[\frac{k^{5}\eta^{2}v^{2}}{4}(1+\eta kR)^{-3}+\eta k^{5}v^{2}(1+\eta kR)^{-2}-8\alpha k^{5}v(1+\eta kR)^{-1/2}\\ -\alpha\eta vk^{5}(1+\eta kR)^{-3/2}+2\alpha^{2}k^{5}]=0\end{split}$ (38)
@@
$\begin{split}k\pi r_{c}[-\frac{3}{4}\eta^{3}v^{2}-2\eta^{2}v^{2}+4\alpha v\eta+\frac{3}{2}\alpha\eta^{2}v]=\alpha\eta v+8\alpha v-\frac{\eta^{2}v^{2}}{4}-\eta v^{2}-2\alpha^{2}\end{split}$ (39)
@@
Thus, we arrive at an expression for the stabilized modulus for a bulk quartic scalar potential. It is indeed easy to see from Eq.(40) that if we were to work with an inverted quartic well ($\zeta\rightarrow-\zeta$), $r_{c}$ acquires a negative value which is unphysical. Hence, just like the standard GW case, we do not get modulus stabilization for the inverted case.
@@
4.4 Bulk double-well potential
@@
We use Eq.(3.1) to find the OR solution for $\Phi$ which takes the form
@@
Figure 3: Effective radion potential (double well) for parameter values $b=0.2,v=0.01,k=10,\alpha=10^{-5}$ with $\lambda=0.001$ (left) and $\lambda=-0.001$ (right)
@@
What we get is that the effective potential has a global minimum for a double well bulk potential i.e. the modulus gets stabilized to a non-trivial non-zero value, which for our benchmark choice of parameters in Fig. 3 turns out to be 0.31. As a result, $k\pi r_{c}\approx 31$ which very well solves the hierarchy problem. However, for the inverted double well, the modulus gets destabilized (no minimum).
4.5 Bazeia-Furtado-Gomes type bulk potential
Bazeia, Furtado and Gomes (BFG); in Ref.DBazeia_2004, introduced a set of bulk scalar potentials which are generated from a class of superpotentials of the form
@@ 4.5 Bazeia-Furtado-Gomes type bulk potential
$V_{b,p}(\Phi)=\frac{1}{8}\left(\frac{dW_{p}}{d\Phi}\right)^{2}-\frac{1}{3}W^{2}_{p}$ (48)
@@
The speciality of these potentials is that the coupled gravity-scalar system gives rise to thick brane solutions. What we do here is take such forms of bulk potentials and plug them in the RS-I background with proper brane potentials - hence solving for the scalar profile and backreacted metric using the superpotential approach of Ref. Freedman (see Appendix A). The next step is to calculate the effective radion potential and check the behaviour for the entire class of superpotentials. A final question is if we can establish a relation between the value of the stabilized modulus ($r_{c}$) [if it gets stabilized] and the parameter p, which is addressed in Appendix B.
Following the steps as in Appendix A, we obtain the following solutions for the scalar profile and the warp factor DBazeia_2004:
@@
$\begin{split}\noindent A_{p}(y)\!\!=\!\!-\frac{1}{3}\frac{p}{2p+1}\tanh^{2p}\left(\frac{y}{p}\right)-\frac{2}{3}\left(\frac{p^{2}}{2p-1}-\frac{p^{2}}{2p+1}\right)\biggl\{\ln[\cosh\left(\frac{y}{p}\right)\biggr]-\sum_{n=1}^{p-1}\frac{1}{2n}\tanh^{2n}\left(\frac{y}{p}\right)\biggr\}\end{split}$ (50)
One should notice that, unlike all the cases till now, this is an exact solution involving the backreaction. The additional inputs that we need to give are the brane potentials which are precisely determined by Eqs. (95) and (96). Since the scalar field vanishes at the hidden brane and so does the superpotential evaluated there, $\lambda_{h}$ does not contribute to the radion potential. On the other hand, the $\lambda_{v}$ contribution to the radion potential is precisely given by $W(\Phi_{v}).e^{{A_{p}(vis)}}$. Hence, putting in the scalar field solution and modified warp factor into the combined gravity-scalar action, integrating over the extra dimension and adding the contribution due to brane potentials; we get the effective radion potential  [^5] The Ricci scalar in the BFG convention is given by $R=8A"+20A^{\prime 2}$. This is incorporated in the numerics.. In this scenario as well, the calculations are quite cumbersome to do analytically, especially when our sole purpose is to determine the presence of modulus stabilization. So, we resort to a numerical approach.
@@
4.6 Mishra-Randall ansatz
Recently, Mishra and Randall Randall_2023 explored the effect of including bulk interaction terms for the GW scalar on RS cosmology. In particular, they were interested in the form of the beta function ($\beta=1/T$), where T is the
Figure 6: GW bulk potentials [($\epsilon_{2}<0,\epsilon_{3}<0$);($\epsilon_{2}<0,\epsilon_{3}>0$); ($\epsilon_{2}>0,\epsilon_{3}<0$); ($\epsilon_{2}>0,\epsilon_{3}>0$)]: from left to right
Figure 7: Radion potentials [($\epsilon_{2}<0,\epsilon_{3}<0$);($\epsilon_{2}<0,\epsilon_{3}>0$); ($\epsilon_{2}>0,\epsilon_{3}<0$); ($\epsilon_{2}>0,\epsilon_{3}>0$)]: from left to right. $10^{67}V(\varphi)$ labels the vertical axis and $\varphi$ labels the horizontal axis.
@@ 4.6 Mishra-Randall ansatz
$\lambda_{hid}(\Phi)=\beta_{hid}(\Phi-v_{hid})^{2}\>,\;\>\beta_{hid}\to\infty\>,\\ \>\lambda_{vis}(\Phi)=2\alpha_{vis}\Phi$ (52)
@@
with $0\leq r\leq r_{vis}$. Plugging everything into the GW action and integrating over the fifth dimension gives the following form for the radion potential Randall_2023
$\begin{split}V_{GW}(\varphi)=\varphi^{4}(A+B\frac{\lambda\varphi^{\epsilon_{2}}}{1-\lambda\varphi^{\epsilon_{2}}}+C(\frac{\lambda\varphi^{\epsilon_{2}}}{1-\lambda\varphi^{\epsilon_{2}}})^{3}-D\log(1-\lambda\varphi^{\epsilon_{2}}))\end{split}$ (54)
@@
$\begin{split}\qquad\varphi=e^{-r_{vis}},\>\lambda=\frac{v_{\text{hid}}\epsilon_{32}}{1+v_{\text{hid}}\epsilon_{32}},\>\epsilon_{32}=\frac{\epsilon_{3}}{\epsilon_{2}},\>A=\frac{1}{768}\epsilon_{3}\alpha_{vis}^{2}-\frac{\lambda}{1-\lambda}+\log(1-\lambda),\\ \>B=-\frac{1}{32}\epsilon_{2}\alpha_{vis}^{2}-\frac{1}{2}\frac{\epsilon_{2}}{\epsilon_{3}}\alpha_{vis},\>C=\frac{1}{6}\frac{\epsilon_{2}^{3}}{\epsilon_{3}^{2}},\>D=\frac{1}{2}\frac{\epsilon_{2}}{\epsilon_{3}}\alpha_{vis}\end{split}$ (55)
We work with a specific choice of parameters, which was part of the benchmark choice in Ref.Randall_2023 as well: $\epsilon_{2}=-1/25,\epsilon_{3}=-1/90,\alpha_{ir}=5/2,v_{uv}=1/5$. The resulting radion potential does have a non-trivial minimum. We now work with the same magnitudes for $\epsilon_{2}$ and $\epsilon_{3}$ but with different sign combinations: namely one of $\epsilon_{2}$ or $\epsilon_{3}$ being positive or both being positive. In the first case, we get a monotonically increasing radion potential with a minimum at $\varphi=0$, corresponding to $r_{vis}\to\infty$ but in the second case, the radion potential has a non-trivial minimum. Thus, we see that the essential condition for radion stabilization in this scenario is that the quadratic and cubic coefficients in the bulk scalar potential must share the same signature, or simply, $\epsilon_{32}>0$.
4.7 General comments
Let us summarize all the cases that we discussed till now. (a) No bulk interaction: only a positive mass-squared bulk scalar ensures radion stabilization. (b) Bulk cubic interaction: only a negative cubic coefficient in the bulk potential stabilizes the radion. (c) Bulk quartic interaction: only a positive quartic coefficient in the bulk potential results in modulus stabilization. (d) Bulk double-well interaction: a bulk double-well potential stabilizes the radion but an inverted well does not. (e) BFG type bulk potential: ensures radion stabilization for any BFG superpotential. (f) Mishra-Randall (MR) ansatz: modulus stabilization happens only when quadratic and cubic coefficients share the same signature. Except for Sec.4.1 and the BFG case, all the cases that we discussed had a GW-type Dirichlet boundary condition on the hidden brane and a linear brane potential on the visible brane, ensuring a consistent comparison. The equivalent analysis for Sec.4.1 with CMS boundary conditions can be found in Sec.5.2. The BFG case was carried out only to demonstrate a non-trivial backreacted geometry. It is known that in the superpotential method, one usually fixes the values of the bulk scalar on the Planck and TeV brane, and this stabilizes the radion - thus it might appear that there is no point in studying such a scenario. But our situation was a bit modified - the scalar field value on the Planck brane was fixed (which serves as an initial condition for the scalar solution), but on the visible brane was not. Instead upon plugging in the scalar profile in the 5D action, we performed dimensional reduction as previously to get the radion potential. The fact that we still got radion stabilization is thus no longer trivial.
The question now is - what possible conjectures could one draw from these cases? For bulk potentials with a local (and no global) minimum, the radion may or may not be stabilized as per the inverted double-well case and the different scenarios of MR ansatz. The same goes for bulk potentials unbounded above and below - both for the cubic case and the variants of MR ansatz, selective modulus stabilization was observed. Thus, though it might be worthwhile to explore general restrictions on the parameters under these conditions (case-specific restrictions were already discussed), there is nothing one could conjecture purely from the qualitative nature of the bulk potential.
A consistent observation is that anywhere the bulk potential had a global minimum, it resulted in radion stabilization, hinting at a generalization of this statement. This was seen to be true for the BFG potentials as well since they have a global minimum, though under different boundary conditions. Again, all the cases where the bulk potential had a global maximum but no global minimum failed to stabilize the radion - can this inference be generalized? A subset of these cases is infact those unstable bulk scalars which violate the BF bound - is there a further connection to scalar field stability? These are the conjectures that rely entirely on the qualitative nature of the bulk scalar potential and are addressed in the next section.
5 Addressing the correspondence: consequences of an extremum in the bulk potential
5.1 Global minimum for $V_{b}(\Phi)$